Cross each allele separately to get the final genotype:
AA x Aa = AA, aa, Aa, Aa .: 1/2
Bb x BB = BB, BB, Bb, Bb .: 1/2
cc x CC = Cc .: 1
Dd x dd = DD, dd, Dd, Dd .: 1/2
Ee x Ee = EE, Ee, Ee, ee .: 1/4
FF x ff = Ff, Ff, Ff, Ff .: 1
Multiply all probable fractions:
1/2 x 1/2 x 1 x 1/2 x 1/4 x 1 = 1/64 chance of that specific genotype.
3/5 probability of a white ball being drawn.
What are the first ten multiples of4?
the multiples of 4 are 4,8,12,16,20,24,28,32,36,40,44,48,52,56,60 those are the multiples of 4
What are the possible outcomes when a die is rolled?
If a normal die is rolled once, the outcomes are {1, 2, 3, 4, 5, 6}.
What is the probability that a gaussian probability density function will take a value less than -2?
Assuming that you are considering a N(0,1) Gaussian distribution, the answer is approximately 1 in 20. The 0.95%ile (two tailed) occurs at -1.96 and 1.96.
50-50
Word 1) 'math' has one vowel letter among a total of 4 letters. The probability of
randomly selecting the vowel letter 'a' is P(v) = 1/4.
Word 2) 'jokes' has two vowel letters among a total of 5 letters. The probability of
randomly selecting a vowel letter is P(v) = 2/5.
The probability of randomly selecting a vowel letter from the first word and a vowel
letter from the second word is: P(v1,v2) = 1/4 (2/5) = 2/20 = 1/10 = 0.10 = 10.0%
How do you find the probability of a number?
Here's a website that might help you with finding the probability of a number:
http://www.mathgoodies.com/lessons/vol6/intro_probability.html
Hope this will help you!
3
When did they discover probability?
A gambler's dispute in 1654 led to the creation of a mathematical theory of probability by two famous French mathematicians, Blaise Pascal and Pierre de Fermat. Antoine Gombaud, Chevalier de Méré, a French nobleman with an interest in gaming and gambling questions, called Pascal's attention to an apparent contradiction concerning a popular dice game. The game consisted in throwing a pair of dice 24 times; the problem was to decide whether or not to bet even money on the occurrence of at least one "double six" during the 24 throws. A seemingly well-established gambling rule led de Méré to believe that betting on a double six in 24 throws would be profitable, but his own calculations indicated just the opposite. This problem and others posed by de Méré led to an exchange of letters between Pascal and Fermat in which the fundamental principles of probability theory were formulated for the first time. Although a few special problems on games of chance had been solved by some Italian mathematicians in the 15th and 16th centuries, no general theory was developed before this famous correspondence. The Dutch scientist Christian Huygens, a teacher of Leibniz, learned of this correspondence and shortly thereafter (in 1657) published the first book on probability; entitled De Ratiociniis in Ludo Aleae, it was a treatise on problems associated with gambling. Because of the inherent appeal of games of chance, probability theory soon became popular, and the subject developed rapidly during the 18th century. The major contributors during this period were Jakob Bernoulli (1654-1705) and Abraham de Moivre (1667-1754). In 1812 Pierre de Laplace (1749-1827) introduced a host of new ideas and mathematical techniques in his book, Théorie Analytique des Probabilités. Before Laplace, probability theory was solely concerned with developing a mathematical analysis of games of chance. Laplace applied probabilistic ideas to many scientific and practical problems. The theory of errors, actuarial mathematics, and statistical mechanics are examples of some of the important applications of probability theory developed in the l9th century. Like so many other branches of mathematics, the development of probability theory has been stimulated by the variety of its applications. Conversely, each advance in the theory has enlarged the scope of its influence. Mathematical statistics is one important branch of applied probability; other applications occur in such widely different fields as genetics, psychology, economics, and engineering. Many workers have contributed to the theory since Laplace's time; among the most important are Chebyshev, Markov, von Mises, and Kolmogorov. that is the history!!!!!!
Since there are 73 Women, and 18 men, there are 91 people all together.
The probability of the student being a man is 18 men out of the 91 total people. So, 18/91 or .1978.
If you have 5 colors on a spinner what is the probability it would land on red?
Assuming that the colors are balanced, the probability is 1 in 5.
Probability of heads in a coin toss?
1/2 Because there is one side with heads (1/) and the total sides (/2)
Time-Probability Graph
If Two dice are rolled Find the probability of getting a sum greater than 8?
The first die can be from 2 to 6, it can not be one. So, the first die has a 5/6 chance. The second die must be a 2 if 6 is on the first die, 3 if 5 is on the first die, etc. so it has a 1/6 chance. Since the two events must occur ("and" situation) and they are independent:
(1/6) * 5/6 = 5/36.
A second way of thinking about this is, that each die can have values from 1 to 6, so you can make a x-y table (1-6 columns and 1-6 rows) and see 36 combinations, of which only 5 will add to 8:
(2,6) (6,2) (3,5) (5,3) (4,4)
so the answer is 5/36.
We know that there are 36 different possible dice rolls.
Since the dice must be greater than 8, we have the following combinations:
3-6 x2
4-5 x2
4-6 x2
5-5
5-6 x2
6-6
Which is 10, or 5/18.
If you looking for > or = to 8, we must add:
2-6 x2
3-5 x2
4-4 to this list.
Which is 15 => 5/12
The probability of getting all 10 questions right is, P(10) =
(1/2)10 =
0.0009765...
9 questions right, P(9) =
10C9 ∙(1/2)10 =
10∙(1/2)10 ~ 0.009765...
8 questions right, P(8) =
10C8 ∙(1/2)10 ~ 45∙(1/2)10 ~ 0.043945...
7 questions right, P(7) =
10C7 ∙(1/2)10 ~ 120∙(1/2)10 ~ 0.1171875...
The probability of passing the test with any of the grade is the sum of all the
above; P(passing)
~ 0.1719 ~ 17.2%
What would be the probability that two straight haired parents would have a curly haired child?
1/4 100% however if one parents hair is even slightly wavy, the child could have curly hair